• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reinhold

    Log entry time 16:49:13 on April 6,2001

    Entry number 59995

    keyword=target status/ shift schedule

    E98-108 mailing list

    Dear Collaborator,

    earlier things looked like we would get beam tonight. The target motion
    had been fixed and the target was being aligned and surveyed.
    Unfortunately later something broke again. That seems to be fixed now.
    Right this minute, the target is being aligned and surveyed again.
    Provided everything will work out fine, we will attempt a cool-down of the
    target early tomorrow morning (Saturday):

    Swing shift Friday and owl shift Saturday are cancelled.
    The Saturday day shift should be available. However, you can take your
    time. Check with the counting house at 10 am.

    Regards,

    Joerg